What is the value of X if 7853-(2853-X)=5500 ?
Answer by macston(5194)

7853-(2853-X)=5000 multiply by -1 to remove parenthesis
7853+(-2853)+(-x)=5000 subtract 5000 from each side
7853-2853-x-5000=0
5000-x-5000=0
-x=0 add x to each side
-x+x=0+x
0=x ANSWER x=0
CHECK:
7853-(2853-x)=5000
7853-(2853-0)=5000
7853-2853=5000
5000=5000